Problem 23-5A Pace Labs, Inc. provides mad cow disease testing for both state and federal governmental agricultural agencies. Because the company’s customers are governmental agencies, prices are strictly regulated. Therefore, Pace Labs must constantly monitor and control its testing costs. Shown below are the standard costs for a typical test. Direct materials (2 test tubes @ $1.46 per tube) $ 2.92 Direct labor (1 hour @ $24 per hour) 24.00 Variable overhead (1 hour @ $6 per hour) 6 Fixed overhead (1 hour @ $10 per hour) 10 Total standard cost per test $42.92 The lab does not maintain an inventory of test tubes. Therefore, the tubes purchased each month are used that month. Actual activity for the month of November 2014, when 1,500 tests were conducted, resulted in the following: Direct materials (3,050 test tubes) $ 4,209 Direct labor (1,600 hours) 36,800 Variable overhead 7,400 Fixed overhead 15,000 Monthly budgeted fixed overhead is $14,000. Revenues for the month were $75,000, and selling and administrative expenses were $5,000. Compute the price and quantity variances for direct materials and direct labor. Materials price variance $ Materials quantity variance $ Labor price variance $ Labor quantity variance $ Compute the total overhead variance. Total Overhead variance $ Prepare an income statement for management. PACE LABS, INC. Income Statement For the Month Ended November 30, 2014 $ $ $
Answer #1
Materials price variance $           244 Favorable
AQ x (AP - SP) = 3050 x ($1.38 - $1.46) = -$244
AR = $4209/3050 = $1.38
Materials quantity variance $             73 Unfavorable
SP x (AQ - SQ) = $1.46 x (3050 - 3000) = $73
SQ = 1500 tests x 2 = 3000 test tubes
Labor price variance $       1,600 Favorable
AH x (AR - SR) = 1600 x ($23 - $24) = -$1600
AR = $36800/1600 = $23
Labor quantity variance $       2,400 Unfavorable
SR x (AH - SH) = $24 x (1600 - 1500) = $2400
SH = 1500 tests x 1 hour = 1500 hours
Total overhead variance $           600 Favorable
Actual overheads incurred = $7400 + $15000 = $22400
Budgeted overheads = (1500 x 1 x $6) + $14000 = $9000 + $14000 = $23000
Total overhead variance = $22400 - $23000 = -$600
PACE LABS, INC.
Income Statement
For the Month Ended November 30, 2014
Service revenue 75000
Cost of service provided (at standard) (1500 x $42.92) 64380
Gross profit (at standard) 10620
Variances
Materials price variance 244 F
Materials quantity variance 73 U
Labor price variance 1600 F
Labor quantity variance 2400 U
Overhead variance 600 F
Total variance 29 U
Gross profit (actual) 10591
Selling and administrative expenses 5000
Net income (loss) 5591